# Subjective-Aligned Dateset and Metric for Text-to-Video Quality Assessment

### v2

@@
Abstract
With the rapid development of generative models, Artificial Intelligence-Generated Contents (AIGC) have exponentially increased in daily lives. Among them, Text-to-Video (T2V) generation has received widespread attention. Though many T2V models have been released for generating high perceptual quality videos, there is still lack of a method to evaluate the quality of these videos quantitatively. To solve this issue, we establish the largest-scale Text-to-Video Quality Assessment DataBase (T2VQA-DB) to date. The dataset is composed of 10,000 videos generated by 9 different T2V models. We also conduct a subjective study to obtain each video’s corresponding mean opinion score. Based on T2VQA-DB, we propose a novel transformer-based model for subjective-aligned Text-to-Video Quality Assessment (T2VQA). The model extracts features from text-video alignment and video fidelity perspectives, then it leverages the ability of a large language model to give the prediction score. Experimental results show that T2VQA outperforms existing T2V metrics and SOTA video quality assessment models. Quantitative analysis indicates that T2VQA is capable of giving subjective-align predictions, validating its effectiveness. The dataset and code will be released at https://github.com/QMME/T2VQA.
Keywords:
Text-to-video dataset Video quality assessment Text-to-video generation
Figure 1: Overview of the proposed T2VQA-DB and T2VQA. T2VQA-DB has the largest scale among existing T2V datasets. T2VQA achieves the SOTA performance in evaluating the quality of text-generated videos.
1 Introduction
@@
Besides, the most used metrics for T2V generation, such as IS [36], FVD [42], and CLIPSim [46], fail to reflect real user preferences. IS uses the Inception Network [40] to generate a distribution that reflects image/video quality and diversity. It has been criticized for its inability to evaluate image/video quality precisely. FVD compares the I3D feature [4] distributions of the generated and natural video pair. The drawback of FVD is that obtaining the reference natural video is usually impractical. CLIPSim takes advantage of CLIP [32] to calculate the similarity between the original text and the generated video content. However, it only considers text-video alignment from the image level, excluding the temporal information and perceptual video quality.
Refer to caption (a) T2VZero [20] Refer to caption (b) Ani-Diff [12] Refer to caption (c) VidRD [11] Refer to caption (d) Tune [56] Refer to caption (e) VF [31] Refer to caption (f) LVDM [14] Refer to caption (g) MS [43] Refer to caption (h) Show-1 [60] Refer to caption (i) LaVie [44] (a)
Refer to caption (j) AnimateDiff [12]. MOS: 62.62. Refer to caption (k) LVDM [14]. MOS: 53.78 Refer to caption (l) LaVie [44]. MOS: 77.57 (b)
Figure 2: Video examples generated by prompt: Sunset on the sea. (a) Overview of video frames generated by 9 models. (b) Videos generated by AnimateDiff [12], LVDM [14], LaVie [44], and their MOSs.
@@
Based on T2VQA-DB, we propose a novel model equipped with multi-modality foundation models for better Text-to-Video Quality Assessment (T2VQA). The model utilizes BLIP [27] and Video Swin Transformer (Swin-T) [30] to extract features from text-video alignment and video fidelity perspectives respectively. The features are fused through a cross-attention module, then they are fed into a frozen Large Language Model (LLM) to regress the predicted score. We train and test T2VQA as well as other VQA models on T2VQA-DB. Experimental results show that T2VQA outperforms existing T2V generation metrics and state-of-the-art VQA models, validating its effectiveness in measuring the perceptual quality of text-generated videos.
@@
  1. 1.
    We establish the T2V dataset with the largest scale to date, named T2VQA-DB. The dataset includes 10,000 text-generated video sequences and their corresponding MOSs gained from 27 subjects.
  2. 2.
@@
    The proposed T2VQA outperforms existing metrics for T2V generation and SOTA VQA models on T2VQA-DB and Sora [3] videos, indicating the effectiveness of T2VQA. Qualitative experiments show that T2VQA can benefit in measuring the performance of T2V generation algorithms, giving it practical application prospects.
2 Related Works
2.1 Text-to-video Dataset
@@
To the best of our knowledge, only a few T2V datasets have been proposed. They mainly have the following two issues: (1) Insufficient scale: Chivileva et al. [5] proposes a dataset with 1,005 videos generated by 5 T2V models following [25]. EvalCrafter [28] builds a dataset using 500 prompts and 7 T2V models, resulting in 3,500 videos in total. Similarly, FETV [29] is composed of 2,476 videos generated by 619 prompts, and 4 T2V models. VBench [16] has a larger scale with in total of $\sim$1.7k prompts and 4 T2V models. Such scales are not sufficient for the training of deep learning-based models, and cannot comprehensively represent current T2V algorithms. (2) Limited human annotation: ITU-standard [17] requires at least 15 human annotators for subjective study. The dataset proposed by Chivileva et al. [5] is the only one that meets the standard with 24 annotators involved. Both EvalCrafter and FETV only have 3 users for annotation. VBench [16] does not specify the number of human annotators explicitly.
2.2 Metrics for T2V Generation
@@
IS [36] and FVD [42] are the two most commonly used metrics for evaluating the quality of generated videos. IS uses the Inception feature to present both image/video quality. FVD measures the distance between the generated video and the natural video. However, both metrics are criticized for poor correlation with human visual perception. CLIPSim [46] measures the text-video alignment by using CLIP [32]. After measuring the similarity between the text and each video frame, it averages them to get the final score. As a result, it only evaluates videos from the image level, losing the information from the temporal domain. As well as it doesn’t consider the video quality.
@@
Though many works targeting the evaluation of text-generated images have been proposed [58, 21, 57, 25, 24, 62], only a few metrics tailored for text-generated video evaluation have been proposed. Among them, ViCLIP [45] is a CLIP-based metric for measuring text-video alignment. Chivileva et al. [5] proposes an ensemble video quality metric that integrates text similarity and naturalness. EvalCrafter [28] and VBench [16] build benchmarks to evaluate text-generated video from 18 and 16 objective metrics respectively. FETV [29] and T2V-Score [55] both propose separate metrics for text-video alignment and video quality, without an overall perceptual score for text-generated videos. There is still lack of a simple and effective metric to evaluate the quality of text-generated videos accurately.
@@
Besides the aforementioned metrics, VQA models can also be used for the evaluation of text-generated videos. BVQA [23] transfers knowledge from Image Quality Assessment (IQA) databases and then trains on the target VQA database. SimpleVQA [39] extracts spatial and motion features to regress to the final score. FAST-VQA [47] proposes “fragments” as a novel sampling strategy and the Fragment Attention Network (FANet) to accommodate fragments as inputs.  [9, 8, 22, 63] are works for the evaluation of enhanced videos and digital humans. DOVER [50] proposes to view the quality assessment problem from the technical perspective and the aesthetic perspective, while BVQI [48] and MaxVQA [49] integrates text prompts (e.g., good, bad) into VQA. With the development of LLMs and Multi-modal Large Language Models (MLLM), researchers have started to leverage the advantages of MLLMs to solve VQA problems, as they have trained on massive data. Q-Bench [51] proves that MLLMs can address preliminary low-level visual tasks. Q-Align [53] proposes to train MLLMs using text-defined levels (e.g., fine, poor, excellent) and achieves SOTA results on both IQA and VQA tasks. Though many VQA models have been proposed, they are originally designed for natural videos and do not consider text-video alignment. Therefore, their performance will deteriorate when evaluating generated videos.
Table 1: Summary of T2VQA-DB and existing T2V datasets. [Keys: Bold: the best].
Name Videos Prompts Models Annotators Score
Chivileva’s [5] 1,005 201 5 24 MOS
EvalCrafter [28] 3,500 500 7 3 MOS
VBench [16] 6,984 1,746 4 - Preference
FETV [29] 2,476 619 4 3 MOS
T2VQA-DB (ours) 10,000 1,000 9 27 MOS
2.3 Text-to-video Generation
@@
T2V generation refers to a form of conditional video generation, where text descriptions are used as conditioning inputs to generate high-fidelity videos. A common practice is to extend pre-trained Text-to-Image (T2I) models with temporal modules. CogVideo [15] is based on CogView2 [7] and proposes a multi-frame-rate hierarchical training strategy to better align text-video clips. Make-a-video [37] adds effective spatial-temporal modules on a diffusion-based T2I model (i.e., DALLE-2 [33]). VideoFusion [31] also leverages the DALLE-2 and presents a decomposed diffusion process. Text2Video-Zero [20], Tune-A-Video [56], AnimateDiff [12], Video LDM [2], MagicVideo[66], LVDM [14], ModelScope [43], and VidRD [11] are models that inherit the success of Stable Diffusion (SD)  [35] for video generation. Show-1 [60] integrates both pixel-based and latent-based text-to-Video Diffusion Models (VDMs). LaVie [44] extends the original transformer block in SD to a spatio-temporal transformer. Recently, OpenAI releases Sora [3], a T2V model that is capable of generating 60s high-fidelity videos, considered as a game changer in T2V generation.
3 Subjective-Aligned Text-to-Video Dataset
@@
3.1 Prompt Selection
To guarantee the diversity of the dataset, the prompts used for T2V generation should cover as many aspects as possible. Following [58], we use the same graph-based algorithm from [38] for prompt selection. We first randomly sample 1 million prompts from WebVid-10M [1], which contains 10 million video-text pairs scraped from the stock footage sites. Each prompt is encoded to a vector representation by Sentence-BERT [34]. The graph-based algorithm integrates them into $k$ groups according to cosine distance. $k$ is a hyper-parameter and we set $k=100$, resulting roughly 10,000 prompts in each group. Finally, we randomly sample 10 prompts in each group, forming the 1,000 prompts in T2VQA-DB. Fig. 2(a) shows the word cloud of the collected prompts.
(a) Word cloud of selected prompts.
(b) The raw/Z-score MOS probability distribution.
(c) The MOSz distributions of 10 T2V models.
Figure 3: (a) The word cloud of the prompts used in T2VQA-DB. (b) The distribution of the raw/Z-score MOS. (c) Z-score MOS distributions of 10 T2V models. The model IDs represent sequentially Text2Video-Zero, Tune-a-video${}^{(1)}$, VidRD, ModelScope, VideoFusion, LVDM, Show-1, Tune-a-video${}^{(2)}$, LaVie.
3.2 Video Generation
We use in total 9 different models for video generation, including Text2Video-Zero [20], AnimateDiff [12], Tune-a-video [56], VidRD [11], VideoFusion [31], ModelScope [43], LVDM [14], Show-1 [60], and LaVie [44]. For Tune-a-video, we utilize two different pre-trained weights, resulting in a total of 10 models for generation. Compared to other T2V datasets, we utilize current advanced T2V generation models as much as possible, making T2VQA-DB more representative. Since the default resolution, video length, and frame rate are different in each model, we unify the video format as $512\times 512$, 16 frames, and 4fps. We end up generating 10,000 text-generated videos.
3.3 Subjective Study
@@
$MOSz_{i}=\frac{1}{N}\sum_{j=1}^{N}Res(\frac{r_{ij}-\mu_{j}}{\sigma_{j}}),$ (1)
where $i$ and $j$ refer to the index of videos and subjects. $r$ is the raw score, and $\mu_{j}=\frac{1}{M}\sum_{i=1}^{M}r_{ij}$, $\sigma_{j}=\sqrt{\frac{1}{M-1}\sum_{i=1}^{M}(r_{ij}-\mu_{j})^{2}}$. $M$ is the number of videos scored by each subject. $N$ is the number of scores on one video. $Res(\cdot)$ is the rescaling function, converting the distribution of Z-scores into a mean of 50 and a standard deviation of 16.6. Fig. 2(b) shows the distributions of raw MOS and Z-score MOS.
Figure 4: (a) Comparison of models performance on different prompt types. (b) Performance of prompt types in different models.
3.4 Dataset Analysis
To have a better understanding of the characteristics of videos in T2VQA-DB, we conduct comprehensive experiments and analysis. We first investigate each model’s performance in T2VQA-DB. The visualization of the Z-score MOS distributions of 10 models is shown in Fig. 2(c). LaVie has the highest average MOS of 66.9, while two Tune-a-Video models have the lowest of 39.1 and 39.9. The reason for the poor performance of Tune-a-Video is mainly the low inter-frame consistency, as shown in Fig. 6(c).
Based on the prompt contents, we categorize the collected prompts into 6 categories, including nature, human, artificial, animal, object, and abstract. The ones that cannot be categorized into the 6 classes are labeled as “others”. After classification, we have 327 prompts for nature, 119 for human, 196 for artificial, 121 for animal, 66 for object, 135 for abstract, and 36 for others. Subsequently, we compare the models’ performance over different prompt types. As shown in Fig. 3(a), LaVie outperforms the other models on all types of prompts. Tune-a-video has the worst performance, which is consistent with the analysis in Fig. 2(c). LVDM has the second-worst performance except in the “others” type. The performance of the rest models has negligible differences. Fig. 3(b) shows prompts classified as human have the worst performance in all models. The reason could be that human faces and actions require more sophisticated modeling compared with other categories. Some models show a preference for object and nature types of prompts to a small extent. The models do not show obvious preferences for other types of prompts.
4 Subjective-Aligned Text-to-video Metric

### v3

# Subjective-Aligned Dataset and Metric for Text-to-Video Quality Assessment

@@
Abstract.
With the rapid development of generative models, AI-Generated Content (AIGC) has exponentially increased in daily lives. Among them, Text-to-Video (T2V) generation has received widespread attention. Though many T2V models have been released for generating high perceptual quality videos, there is still lack of a method to evaluate the quality of these videos quantitatively. To solve this issue, we establish the largest-scale Text-to-Video Quality Assessment DataBase (T2VQA-DB) to date. The dataset is composed of 10,000 videos generated by 9 different T2V models, along with each video’s corresponding mean opinion score. Based on T2VQA-DB, we propose a novel transformer-based model for subjective-aligned Text-to-Video Quality Assessment (T2VQA). The model extracts features from text-video alignment and video fidelity perspectives, then it leverages the ability of a large language model to give the prediction score. Experimental results show that T2VQA outperforms existing T2V metrics and SOTA video quality assessment models. Quantitative analysis indicates that T2VQA is capable of giving subjective-align predictions, validating its effectiveness. The dataset and code will be released at https://github.com/QMME/T2VQA.
Text-to-video dataset and Video quality assessment and Text-to-video generation
Text-to-video dataset, Video quality assessment , Text-to-video generation
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Proceedings of the 32st ACM International Conference on Multimedia; October 28–November 1, 2024; Melbourne, Australiaisbn: 978-1-4503-XXXX-X/18/06submissionid: 678ccs: Computing methodologies Modeling methodologies
Figure 1. Overview of the proposed T2VQA-DB and T2VQA. T2VQA-DB has the largest scale among existing T2V datasets. T2VQA achieves the SOTA performance in evaluating the quality of text-generated videos.
1. Introduction
(a) t
(b) t
(c) t
(d) t
(a)(b)
Figure 2. Video examples generated by prompt: Sunset on the sea. (a) Overview of video frames generated by 9 models. (b) Videos generated by AnimateDiff (Guo et al., 2023), LVDM (He et al., 2022), LaVie (Wang et al., 2023a), and their MOSs.
@@ 1. Introduction
Besides, the most used metrics for T2V generation, such as IS (Salimans et al., 2016), FVD (Unterthiner et al., 2018), and CLIPSim (Wu et al., 2021), fail to reflect real user preferences. IS uses the Inception Network (Szegedy et al., 2016) to generate a distribution that reflects image/video quality and diversity. It has been criticized for its inability to evaluate image/video quality precisely. FVD compares the I3D feature (Carreira and Zisserman, 2017) distributions of the generated and natural video pair. The drawback of FVD is that obtaining the reference natural video is usually impractical. CLIPSim takes advantage of CLIP (Radford et al., 2021) to calculate the similarity between the original text and the generated video content. However, it only considers text-video alignment from the image level, excluding the temporal information and perceptual video quality.
@@
Based on T2VQA-DB, we propose a novel model equipped with multi-modality foundation models for better Text-to-Video Quality Assessment (T2VQA). The model utilizes BLIP (Li et al., 2022a) and Video Swin Transformer (Swin-T) (Liu et al., 2022) to extract features from text-video alignment and video fidelity perspectives respectively. The features are fused through a cross-attention module, then they are fed into a frozen Large Language Model (LLM) to regress the predicted score. We train and test T2VQA as well as other VQA models on T2VQA-DB. Experimental results show that T2VQA outperforms existing T2V generation metrics and state-of-the-art VQA models, validating its effectiveness in measuring the perceptual quality of text-generated videos.
@@
  1. (1)
    We establish the T2V dataset with the largest scale to date, named T2VQA-DB, which includes 10,000 text-generated video sequences and their corresponding MOSs gained from 27 subjects.
  2. (2)
@@
    The proposed T2VQA outperforms existing T2V generation metrics and SOTA VQA models on T2VQA-DB and Sora (Brooks et al., 2024) videos, indicating the effectiveness of T2VQA. Qualitative experiments show that T2VQA can benefit in measuring the performance of T2V generation algorithms, giving it practical application prospects.
2. Related Works
2.1. Text-to-video Dataset
@@
To the best of our knowledge, only a few T2V datasets have been proposed. They mainly have the following two issues: (1) Insufficient scale: Chivileva et al(Chivileva et al., 2023) proposes a dataset with 1,005 videos generated by 5 T2V models following (Li et al., 2023b). EvalCrafter (Liu et al., 2023) builds a dataset using 500 prompts and 7 T2V models, resulting in 3,500 videos in total. Similarly, FETV (Liu et al., 2024) is composed of 2,476 videos generated by 619 prompts, and 4 T2V models. VBench (Huang et al., 2024) has a larger scale with in total of $\sim$1.7k prompts and 4 T2V models. Such scales are not sufficient for the training of deep learning-based models, and cannot comprehensively represent current T2V algorithms. (2) Limited human annotation: ITU-standard (ITU-R, 2002) requires at least 15 human annotators for subjective study. The dataset proposed by Chivileva et al(Chivileva et al., 2023) is the only one that meets the standard with 24 annotators involved. Both EvalCrafter and FETV only have 3 users for annotation. VBench (Huang et al., 2024) does not specify the number of human annotators explicitly.
Table 1. Summary of T2VQA-DB and existing T2V datasets. [Keys: Bold: the best].
Name Videos Prompts Models Annotators
Chivileva’s (Chivileva et al., 2023) 1,005 201 5 24
EvalCrafter (Liu et al., 2023) 3,500 500 7 3
VBench (Huang et al., 2024) 6,984 1,746 4 -
FETV (Liu et al., 2024) 2,476 619 4 3
T2VQA-DB (ours) 10,000 1,000 9 27
2.2. Metrics for T2V Generation
@@
IS (Salimans et al., 2016) and FVD (Unterthiner et al., 2018) are the two most commonly used metrics for evaluating the quality of generated videos. IS uses the Inception feature to present both image/video quality. FVD measures the distance between the generated video and the natural video. However, both metrics are criticized for poor correlation with human visual perception. CLIPSim (Wu et al., 2021) measures the text-video alignment by using CLIP (Radford et al., 2021). After measuring the similarity between the text and each video frame, it averages them to get the final score. As a result, it only evaluates videos from the image level, losing the information from the temporal domain. As well as it doesn’t consider the video quality.
@@
Though many works targeting the evaluation of text-generated images have been proposed (Xu et al., 2024; Kirstain et al., 2024; Wu et al., 2023c; Li et al., 2023b, 2024; Zhang et al., 2023a), only a few metrics tailored for text-generated video evaluation have been proposed. Among them, ViCLIP (Wang et al., 2023b) is a CLIP-based metric for measuring text-video alignment. Chivileva et al(Chivileva et al., 2023) proposes an ensemble video quality metric that integrates text similarity and naturalness. EvalCrafter (Liu et al., 2023) and VBench (Huang et al., 2024) build benchmarks to evaluate text-generated video from 18 and 16 objective metrics respectively. FETV (Liu et al., 2024) and T2V-Score (Wu et al., 2024a) both propose separate metrics for text-video alignment and video quality, without an overall perceptual score for text-generated videos. There is still lack of a simple and effective metric to evaluate the quality of text-generated videos accurately.
@@
Besides the aforementioned metrics, VQA models can also be used for the evaluation of text-generated videos. BVQA (Li et al., 2022b) transfers knowledge from Image Quality Assessment (IQA) databases and then trains on the target VQA database. SimpleVQA (Sun et al., 2022) extracts spatial and motion features to regress to the final score. FAST-VQA (Wu et al., 2022) proposes “fragments” as a novel sampling strategy and the Fragment Attention Network (FANet) to accommodate fragments as inputs.  (Gao et al., 2023; Dong et al., 2023; Kou et al., 2023; Zhang et al., 2023b) are works for the evaluation of enhanced videos and digital humans. DOVER (Wu et al., 2023e) proposes to view the quality assessment problem from the technical perspective and the aesthetic perspective, while BVQI (Wu et al., 2023b) and MaxVQA (Wu et al., 2023d) integrates text prompts (e.g., good, bad) into VQA. With the development of LLMs and Multi-modal Large Language Models (MLLM), researchers have started to leverage the advantages of MLLMs to solve VQA problems, as they have trained on massive data. Q-Bench (Wu et al., 2023g) proves that MLLMs can address preliminary low-level visual tasks. Q-Align (Wu et al., 2023f) proposes to train MLLMs using text-defined levels (e.g., fine, poor, excellent) and achieves SOTA results on both IQA and VQA tasks. Though many VQA models have been proposed, they are originally designed for natural videos and do not consider text-video alignment.
2.3. Text-to-video Generation
@@
T2V generation refers to a form of conditional video generation, where text descriptions are used as conditioning inputs to generate high-fidelity videos. A common practice is to extend pre-trained Text-to-Image (T2I) models with temporal modules. CogVideo (Hong et al., 2022) is based on CogView2 (Ding et al., 2022) and proposes a multi-frame-rate hierarchical training strategy to better align text-video clips. Make-a-video (Singer et al., 2022) adds effective spatial-temporal modules on a diffusion-based T2I model (i.e., DALLE-2 (Ramesh et al., 2022)). VideoFusion (Luo et al., 2023) also leverages the DALLE-2 and presents a decomposed diffusion process. LVDM (He et al., 2022), Text2Video-Zero (Khachatryan et al., 2023), Tune-A-Video (Wu et al., 2023a), AnimateDiff (Guo et al., 2023), Video LDM (Blattmann et al., 2023), MagicVideo(Zhou et al., 2022), ModelScope (Wang et al., 2023c), and VidRD (Gu et al., 2023) are models that inherit the success of Stable Diffusion (SD)  (Rombach et al., 2022) for video generation. Show-1 (Zhang et al., 2023d) integrates both pixel-based and latent-based text-to-Video Diffusion Models (VDMs). LaVie (Wang et al., 2023a) extends the original transformer block in SD to a spatio-temporal transformer. Recently, OpenAI releases Sora (Brooks et al., 2024), a T2V model that is capable of generating 60s high-fidelity videos, considered as a game changer in T2V generation.
3. Subjective-Aligned Text-to-Video Dataset
(a) Word cloud of selected prompts.
(b) The raw/Z-score MOS distribution.
(c) The MOSz distributions of 10 T2V models.
Figure 3. (a) The word cloud of the prompts used in T2VQA-DB. (b) The distribution of the raw/Z-score MOS. (c) Z-score MOS distributions of 10 T2V models. The model IDs represent sequentially Text2Video-Zero, Tune-a-video(1), VidRD, ModelScope, VideoFusion, LVDM, Show-1, Tune-a-video(2), LaVie.
@@ 3. Subjective-Aligned Text-to-Video Dataset
3.1. Prompt Selection
To guarantee the diversity of the dataset, the prompts used for T2V generation should cover as many aspects as possible. Following (Xu et al., 2024), we use the same graph-based algorithm from (Su et al., 2022) for prompt selection. We first randomly sample 1 million prompts from WebVid-10M (Bain et al., 2021), which contains 10 million video-text pairs scraped from the stock footage sites. Each prompt is encoded to a vector representation by Sentence-BERT (Reimers and Gurevych, 2019). The graph-based algorithm integrates them into $k$ groups according to cosine distance. $k$ is a hyper-parameter and we set $k=100$, resulting roughly 10,000 prompts in each group. Finally, we randomly sample 10 prompts in each group, forming the 1,000 prompts in T2VQA-DB. Fig. 3a shows the word cloud of the collected prompts.
3.2. Video Generation
We use in total 9 different models for video generation, including Text2Video-Zero (Khachatryan et al., 2023), AnimateDiff (Guo et al., 2023), Tune-a-video (Wu et al., 2023a), VidRD (Gu et al., 2023), VideoFusion (Luo et al., 2023), ModelScope (Wang et al., 2023c), LVDM (He et al., 2022), Show-1 (Zhang et al., 2023d), and LaVie (Wang et al., 2023a). For Tune-a-video, we utilize two different pre-trained weights, resulting in a total of 10 models for generation. Compared to other T2V datasets, we utilize current advanced T2V generation models as much as possible, making T2VQA-DB more representative. Since the default resolution, video length, and frame rate are different in each model, we unify the video format as $512\times 512$, 16 frames, and 4fps. We end up generating 10,000 text-generated videos.
3.3. Subjective Study
@@
(1) $MOSz_{i}=\frac{1}{N}\sum_{j=1}^{N}Res(\frac{r_{ij}-\mu_{j}}{\sigma_{j}}),$
where $i$ and $j$ refer to the index of videos and subjects. $r$ is the raw score, and $\mu_{j}=\frac{1}{M}\sum_{i=1}^{M}r_{ij}$, $\sigma_{j}=\sqrt{\frac{1}{M-1}\sum_{i=1}^{M}(r_{ij}-\mu_{j})^{2}}$. $M$ is the number of videos scored by each subject. $N$ is the number of scores on one video. $Res(\cdot)$ is the rescaling function, converting the distribution of Z-scores into a mean of 50 and a standard deviation of 16.6. Fig. 3b shows the distributions of raw MOS and Z-score MOS.
Figure 4. Comparison of models performance on different prompt types.
3.4. Dataset Analysis
Figure 5. Overview framework of T2VQA. Features from text-video alignment and video fidelity perspectives are extracted. After a cross-attention based fusion, an LLM is utilized for regression.
We also conduct comprehensive experiments and analysis on T2VQA-DB. We first investigate each model’s performance in T2VQA-DB. The visualization of the Z-score MOS distributions of 10 models is shown in Fig. 3c. LaVie has the highest average MOS of 66.9, while two Tune-a-Video models have the lowest of 39.1 and 39.9. The reason for the poor performance of Tune-a-Video is mainly the low inter-frame consistency, as shown in Fig. 7c.
Based on the prompt contents, we classify the collected prompts into 6 categories, including nature, human, artificial, animal, object, and abstract. The ones that cannot be categorized into the 6 classes are labeled as “others”. After classification, we have 327 prompts for nature, 119 for human, 196 for artificial, 121 for animal, 66 for object, 135 for abstract, and 36 for others. Subsequently, we compare the models’ performance over different prompt types. As shown in Fig. 4, LaVie outperforms the other models on all types of prompts. Tune-a-video has the worst performance, which is consistent with the analysis in Fig. 3c. LVDM has the second-worst performance except in the “others” type. The performance of the rest models has negligible differences. Fig. 4 also shows prompts classified as human have the worst performance in all models. The reason could be that human faces and actions require more sophisticated modeling compared with other categories. Some models show a preference for object and nature types of prompts to a small extent. The models do not show obvious preferences for other types of prompts.
4. Subjective-Aligned Text-to-video Metric